In planning for a new item, a manufacturer assumes that the number of items produced X and the cost of dollars C of producing these items are related by a linear equation. Projections are that 100 items will cost $10,000 to produce and that 300 items will cost $22,000 to produce.Find the equation that relates to C and X.
McGraw Hill 6th edition. CH. 7.3 Problem # 52.
Thank you so very much for your time!!!!!

Answer by stanbon(48574)

Projections are that 100 items will cost $10,000 to produce and that 300 items will cost $22,000 to produce.Find the equation that relates to C and X.
-----------
You have two points: (100,10000) and (300,22000)
The slope is [22000-10000]/[300-100]= 12000/200=60
The form is C=mX+b; you need to find "b".
10000=60(100)+b
10000=6000+b
b=4000
EQUATION:
Cost=60X+4000
Cheers,
Stan H.